# Quantum dispersionless KdV hierarchy revisited

### v2

@@ 3.3. Properties of quantum Hamiltonians
We notice that each $\varphi(\tilde{h}_{n})$ is of conformal degree $n+1$, hence $\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})$ is of conformal degree $n+m+1$, which implies that
$\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})\in TV.$
This proves Theorem 1.3 for $\tilde{h}_{n}$.
@@
Finally, we are to prove $\tilde{h}_{n}=h_{n}$. Due to $\tilde{h}_{0}=h_{0}$, it suffices to verify the recursion relation
@@
Let us prove the validity of the above relation by induction on $n$. The case $n=0$ can be directly verified, and we assume that we have proved this for $n=0,1,\dots,N-1$ for some $N\geq 1$. Now let us consider the case $n=N$. Firstly, it follows from the recursion
@@
=\,= $\frac{2}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{1}{\binom{N+1}{k+1}}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{N-1-k})$
@@ 4. Conclusion
In this paper, we study the relationship between vertex algebras and quantum integrable hierarchies. This idea is quite natural, indeed, Barakat, De Sole, Kac and their collaborators developed a theory to describe Hamiltonian integrable hierarchies via Poisson vertex algebras (see. e.g., [13]), and these algebras, in some cases, can be viewed as a classical limit of vertex algebras. In this sense, the vertex algebras should provide suitable tools to study quantum integrable hierarchies, and we investigate this possibility in this paper via the example of quantum dispersionless KdV hierarchy.
Let us compare the results of the present paper to those obtained in the papers [2, 5]. In the first place, a different deformation quantization is constructed in [2], which is given by the commutator of the standard Moyal product corresponding to the Hamiltonian structure. In terms of differential polynomials, the deformed Hamiltonian structure reads
$[F,G]=\int\sum_{m\geq 1}\frac{\hbar^{m}}{m!}$ $\sum_{\begin{subarray}{c}r_{1},\dots,r_{m}\geq 0\\ s_{1},\dots,s_{m}\geq 0\end{subarray}}\frac{\partial^{m}f}{v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}}}\prod_{i=1}^{m}(-1)^{r_{i}}\eta^{\alpha_{i}\beta_{i}}$
×$\times\sum_{j=1}^{2m-1+\sum r_{i}+\sum s_{i}}C_{j}^{r_{1}+s_{1}+1,\dots,r_{m}+s_{m}+1}\partial_{x}^{j}\left(\frac{\partial^{m}g}{\partial v^{\alpha_{1},r_{1}}\dots\partial v^{\alpha_{m},r_{m}}}\right),$
where $C_{j}^{a_{1},\dots,a_{m}}$ are certain constants, and in particular we have
@@
Therefore, the above bracket can be viewed as a further deformation of the bracket constructed in the present paper.
@@
Despite these differences, our results also share similarity to the known results. Therefore, it makes sense to conjecture that our formulation is equivalent to the formulation presented in [2, 5] in some sense. For example, can we actually recover the results in [5] using results of the present paper? Is the above bracket constructed in [2] related to the bracket presented in this paper by a certain transformation? These problems are still under investigation.

### v3

@@ 3.3. Properties of quantum Hamiltonians
We notice that each $\varphi(\tilde{h}_{n})$ is of conformal weight $n+1$, hence $\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})$ is of conformal weight $n+m+1$, which implies that $C_{n,m}=0$ and for $(m,n)\neq(0,0)$ the constants $B_{n,m}$ also vanish, namely
$\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})\in TV,\quad(m,n)\neq(0,0).$
But for $m=n=0$, we know $\varphi(\tilde{h}_{0})=b$ and it follows that $\varphi(\tilde{h}_{0})_{(0)}=0$, from which we conclude that the above identity also holds true in this case. This proves Theorem 1.3 for $\tilde{h}_{n}$.
@@
Finally, we are to prove $\tilde{h}_{n}=h_{n}$. Since we already have $\tilde{h}_{0}=h_{0}=v$, it suffices to prove the validity of the recursion relation
@@
Let us verify the above relation by induction on $n$. The case $n=0$ can be proved directly and we assume that we have proved this for $n=0,1,\dots,N-1$ for some $N\geq 1$. Now let us consider the case $n=N$. Firstly, it follows from the recursion
@@
$=\,$ $\frac{2}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{1}{\binom{N+1}{k+1}}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{N-1-k})$
@@ 4. Conclusion
In this paper, we study the relationship between vertex algebras and quantum integrable hierarchies. This idea is quite natural, indeed, Barakat, De Sole, Kac and their collaborators developed a theory to describe Hamiltonian integrable hierarchies via Poisson vertex algebras (see, e.g., [14]), and these algebras, in some cases, can be viewed as a classical limit of vertex algebras. In this sense, the vertex algebras should provide suitable tools to study quantum integrable hierarchies, and we investigate this possibility in this paper via the example of quantum dispersionless KdV hierarchy. An explicit deformation quantization of Hamiltonian structures of hydrodynamic type is constructed, and the corresponding quantum Hamiltonians are constructed for this hierarchy. All these constructions are based on the structure theory of the Heisenberg vertex algebra.
Let us compare the results of the present paper to those obtained in the papers [3, 6]. First, we briefly recall their construction of quantum Hamiltonian structures and quantum Hamiltonians for quantum dispersionless KdV hierarchy. The classical Hamiltonian structure (1.1) can be equivalently represented by
(4.1) $\{v(x),v(y)\}=\delta^{\prime}(x-y).$
This Poisson bracket admits a canonical quantization by expanding the field $v(x)$ into Fourier modes
$v(x)=\sum_{k\in\mathbb{Z}}v_{k}e^{ikx},$
and rewriting the Poisson bracket (4.1) into the form
$\{v_{m},v_{n}\}=i\,m\,\delta_{m+n,0}.$
Then, the Fourier modes $v_{n}$ are quantized to creators $\hat{q}_{n}$ for $n\geq 1$ and to annihilators $\hat{p}_{-n}$ for $n\leq-1$, and the mode $v_{0}$ is quantized to 0 for simplicity. These quantum operators satisfy the commutating relation
(4.2) $\hat{p}_{n},\hat{q}_{m}]=\hbar\,n\,\delta_{m,n},\quad[\hat{p}_{n},\hat{p}_{m}]=[\hat{q}_{n},\hat{q}_{m}]=0,\quad n,m\geq 1,$
where $\hbar$ is a formal parameter. An explicit formula of the above quantum Hamiltonian structure in terms of differential polynomials is given in [3], and it reads
$[F,G]=\int\sum_{m\geq 1}\frac{\hbar^{m}}{m!}$ $\sum_{\begin{subarray}{c}r_{1},\dots,r_{m}\geq 0\\ s_{1},\dots,s_{m}\geq 0\end{subarray}}\frac{\partial^{m}f}{v^{(s_{1})}\dots v^{(s_{m})}}(-1)^{\sum r_{i}}$
$\cdot\sum_{j=1}^{2m-1+\sum r_{i}+\sum s_{i}}C_{j}^{r_{1}+s_{1}+1,\dots,r_{m}+s_{m}+1}\partial_{x}^{j}\left(\frac{\partial^{m}g}{\partial v^{(r_{1})}\dots\partial v^{(r_{m})}}\right),$
where $F,G\in\mathcal{F}$ with $f,g\in\mathcal{A}$ a choice of their densities, and $C_{j}^{a_{1},\dots,a_{m}}$ are certain constants. In particular, it is proved that
@@
We compare this bracket to the one we obtained in Theorem 1.1, and conclude that the above bracket can be viewed as a further deformation of the quantum Hamiltonian structure constructed in the present paper.
The quantum Hamiltonians constructed and studied in [3, 6] are also different from those presented in this paper. Denote by $H_{n}^{BR}=\int h_{n}^{BR}$ the quantum Hamiltonians constructed by Buryak and Rossi, then it is given by the following generating series [6, 21]:
$h^{BR}(z)=1+\sum_{n\geq 0}h_{n}^{BR}z^{n+1}=\exp\left(zs(\sqrt{\hbar}z\partial_{x})v(x)\right),$
where the function
$s(t)=\frac{e^{t/2}-e^{-t/2}}{t}=\sum_{n\geq 0}\frac{t^{2n}}{4^{n}(2n+1)!}.$
Here we slightly modify the original choices of densities presented in [3, 6] for a clearer comparison. Indeed, we compare them with the quantum Hamiltonian densities constructed in the present paper, and they read
$h_{0}=h_{0}^{BR}=v;$
$h_{1}=h_{1}^{BR}=\frac{1}{2}v^{2};$
$h_{2}=\frac{1}{6}v^{3}+\frac{\hbar}{12}v_{xx},\quad h_{2}^{BR}=\frac{1}{6}v^{3}+\frac{\hbar}{24}v_{xx};$
$h_{3}=\frac{v^{4}}{24}+\frac{\hbar}{24}\left(v_{x}^{2}+2vv_{xx}\right),\quad h_{3}^{BR}=\frac{v^{4}}{24}+\frac{\hbar}{24}vv_{xx}.$
The choices of quantum densities are different, however, we note that as local functionals, $H_{n}=H_{n}^{BR}$ in the above examples. Actually, this can be checked for more examples, and we conjecture this to hold true for any $n$. This implies a close relation between two different ways of quantizations.
@@
Despite these differences, our results also share similarity to the known results. Besides the conjectural relation $H_{n}=H_{n}^{BR}$, the common eigenvectors of $H_{n}^{BR}$ computed in [6] are also given by Schur polynomials, and further properties of quantum KdV hierarchy studied in [11, 12, 22] seem to be applicable to the present setting. Therefore, it makes sense to conjecture that our formulation is equivalent to the formulation presented in [3, 6] in some sense. For example, can we actually recover the results in [6] using results of the present paper? Is the bracket constructed in [3] related to the bracket presented in this paper by a certain transformation? These problems are still under investigation.